### v2

# Deep learning methods for inverse problems using connections between proximal operators and Hamilton–Jacobi equations

Abstract
Inverse problems are important mathematical problems that seek to recover model parameters from noisy data. Since inverse problems are often ill-posed, they require regularization or incorporation of prior information about the underlying model or unknown variables. Proximal operators, ubiquitous in nonsmooth optimization, are central to this because they provide a flexible and convenient way to encode priors and build efficient iterative algorithms. They have also recently become key to modern machine learning methods, e.g., for plug-and-play methods for learned denoisers and deep neural architectures for learning priors of proximal operators. The latter was developed partly due to recent work characterizing proximal operators of nonconvex priors as subdifferential of convex potentials. In this work, we propose to leverage connections between proximal operators and Hamilton–Jacobi partial differential equations (HJ PDEs) to develop novel deep learning architectures for learning the prior. In contrast to other existing methods, we learn the prior directly without recourse to inverting the prior after training. We present several numerical results that demonstrate the efficiency of the proposed method in high dimensions.
1 Introduction
Inverse problems are ubiquitous mathematical problems that primarily aim at recovering model parameters from noisy data. They arise in many scientific and engineering applications for, e.g., recovering an image from noisy measurements, deblurring, tomographic reconstruction, and compressive sensing [AEOV2023, arridge2019solving, bertero2021introduction, isakov2017inverse]. Since inverse problems are often ill-posed, it is essential to include regularization or prior information about the underlying model or unknown variables. Proximal operators are central to this: they provide a flexible and computationally convenient way to encode priors and to build efficient iterative algorithms (e.g., proximal (sub)gradients and splitting methods). More recently, proximal operators have become key ingredients for state-of-the-art machine learning methods, such as plug-and-play methods that replace explicit regularizers by learned denoisers [hu2023plug, jia2025plug] or deep neural architectures parameterizing proximal maps or their gradients. These developments have made proximal methods practical and powerful computational tools.
Formally, the proximal operator of a proper function $J\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is defined via an observed data $\bm{x}\in\mathbb{R}^{n}$, a parameter $t>0$, and the minimization problem
(1) $S(\bm{x},t)=\min_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J(\bm{y})\right\}.$
The proximal operator $\text{prox}_{tJ}\colon\mathbb{R}^{n}\to\mathbb{R}$ is the set-valued function
(2) $\text{prox}_{tJ}(\bm{x})=\operatorname*{arg\,min}_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J(\bm{y})\right\}.$
Here, $t$ controls the trade-off between the quadratic data-fidelity term and the prior $J$. In practice one often works directly with $\text{prox}_{tJ}$ rather than the prior.
The recent work of Gribonval and Nikolova [gribonval2020characterization] in nonsmooth optimization has extended the characterization of proximal operators with convex priors to those with nonconvex priors, showing they are functions that are subdifferentials of certain convex potentials. These properties, in particular, were used in [fang2024whats] to develop new deep learning methods, called learned proximal networks (LPNs), to learn from data the underlying prior of a proximal operator.
The paper [gribonval2020characterization] did not, however, discuss the well-established, existing connections between proximal operators and Hamilton–Jacobi Partial Differential Equations (HJ PDEs) [chaudhari2018deep, darbon2015convex, darbon2021bayesian, darbon2016algorithms, osher2023hamilton]. To see these connections, consider the following HJ PDE with quadratic Hamiltonian function and whose initial data is the prior $J$:
(3) $\begin{dcases}\frac{\partial S}{\partial t}(\bm{x},t)+\frac{1}{2}\left\|{\nabla_{\bm{x}}S(\bm{x},t)}\right\|_{2}^{2}=0,&\ \bm{x}\in\mathbb{R}^{n}\times(0,+\infty),\\ S(\bm{x},0)=J(\bm{x}),&\ \bm{x}\in\mathbb{R}^{n}.\end{dcases}$
If $J$ is uniformly Lipschitz continuous, then the unique viscosity solution of the HJ PDE is given by Eq. 1. Moreover, at a point of differentiability $\bm{x}$, there holds
(4) $\text{prox}_{tJ}(\bm{x})=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t).$
Finally, the viscosity solution is semiconcave, that it, the function $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t)$ is convex. Thus, when paired with Eq. 4, the function $\text{prox}_{tJ}(\bm{x})$ is obtained from differentiating a convex function. This formally connects proximal operators to HJ PDEs (which we emphasize was previously known and established) and the stronger [^1] To the best our knowledge, this characterization result was unknown in the theory of HJ PDEs. characterization obtained in [gribonval2020characterization].
In this paper, we propose to leverage the theory of viscosity solutions of HJ PDEs to develop novel deep learning methods for learning the prior function $J$ in Eq. 2 directly from data. To describe our approach, suppose the solution $(\bm{x},t)\mapsto S(\bm{x},t)$ to the HJ PDE Eq. 3 is known. (We will consider the case when only samples of it are known later.) This problem was investigated in [barron1999regularity, claudel2011convex, colombo2020initial, esteve2020inverse, misztela2020initial]. In particular, [esteve2020inverse] showed that when $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t)$ is convex, there exists a prior $J$ recovering $S(\bm{x},t)$ exactly. Moreover, there is a natural choice for the prior, obtained by reversing the time in the HJ PDE Eq. 3 and using $(\bm{x},t)\mapsto S(\bm{x},t)$ as the terminal condition. The resulting backward viscosity solution yields the prior $J_{\text{BVS}}\colon\mathbb{R}^{n}\to\mathbb{R}$ which admits the representation formula
(5) $J_{\text{BVS}}(\bm{y})=\sup_{\bm{x}\in\mathbb{R}^{n}}\left\{S(\bm{x},t)-\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}\right\}.$
Here, $J(\bm{y})\geqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, with $J_{\text{BVS}}(\bm{y})=J(\bm{y})$ whenever $\bm{y}=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$, where $\bm{x}$ is a point of differentiability of $\bm{x}\mapsto S(\bm{x},t)$. Moreover,
$\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{BVS}}(\bm{y})\right\}=S(\bm{x},t)\ \text{for every $\bm{x}\in\mathbb{R}^{n}$.}$
Thus the prior $J_{\text{BVS}}$ recovers the function $x\mapsto S(\bm{x},t)$, although in general $\text{prox}_{tJ}$ and $\text{prox}_{tJ_{\text{BVS}}}$ may not agree everywhere. Nonetheless, this provides a principled way to estimate the prior, at least when $S(\bm{x},t)$ is known.
We shall focus in this paper on the case where $\bm{x}\mapsto S(\bm{x},t)$ is unknown but have access to some samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ with $t$ fixed. We propose to learn the prior $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})$ by leveraging the crucial fact that $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is convex, thus enabling approaches based on deep learning and convex neural networks.
Related works: Hamilton–Jacobi PDEs are important to many scientific and engineering applications arising in e.g., optimal control [Bardi1997Optimal, fleming2006controlled, mceneaney2006max, parkinson2018optimal] and physics [Caratheodory1965CalculusI, Caratheodory1967CalculusII], inverse problems for imaging sciences [darbon2015convex, darbon2021bayesian, darbon2019decomposition, darbon2022hamilton], optimal transport [meng2024primal, onken2021ot], game theory [BARRON1984213, Evans1984Differential, ruthotto2020machine], and machine learning [chen2024leveraging, zou2024leveraging]. Recent works focus on developing specialized methods for solving high-dimensional HJ PDEs via e.g., representation formulas or deep learning. These specialized methods leverage certain properties of HJ PDEs, including stochastic aspects and representation formulas [bardi1998hopf, darbon2022hamilton, darbon2016algorithms, mceneaney2006max], to approximate solutions to HJ PDEs more accurately and efficiently than general-purpose methods. See, e.g., [darbon2023neural, darbon2020overcoming, darbon2021some, meng2022sympocnet, park2025neural] for recent works along these lines and [meng2025recent] for a review of the state-of-the-art numerical methods for HJ PDEs.
Deep learning methods have become popular for computing solutions to high-dimensional PDEs and their inverse problems. This is because neural networks can be trained on data to approximate high-dimensional, nonlinear functions using efficient optimization algorithms. They have been used to approximate solutions to PDEs without any discretization with numerical grids, and for this reason they can overcome, or at least mitigate, the curse of dimensionality. There is a fairly comprehensive literature on deep learning methods for solving PDEs in general, e.g., see [beck2020overview, cuomo2022scientific, karniadakis2021physics].
Organization of this paper: We present background information on proximal operators, Hamilton–Jacobi equations, and convex neural networks in Section 2. Next, we discuss recent results concerning the inverse problem for Hamilton–Jacobi equations when the solution is available, and how they relate to proximal operators and learning priors in inverse problems, in Section 3. We present some theoretical results in Section 4, where we study the inverse problem for Hamilton–Jacobi equations when only incomplete information is available about its solution. We suggest via arguments from max-plus algebra theory for Hamilton–Jacobi PDEs how to learn from data the solution to a certain Hamilton-Jacobi–Jacobi terminal value problem, which can then be used as an estimate for learning the prior function in a proximal operator. We present in Section 5 some numerical experiments for learning the initial data of certain Hamilton–Jacobi PDEs using convex neural networks and the theory of inverse Hamilton–Jacobi PDEs. Finally, we summarize our results in Section 6.
2 Background
We present here some background on proximal operators, HJ PDEs, connections between them, and convex neural networks. For comprehensive references, we refer the reader to [cannarsa2004semiconcave, evans2022partial, rockafellar2009variational].
2.1 Proximal operators
Let $J\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ denote a proper function (i.e., $J(\bm{x})<+\infty$ for some $\bm{x}\in\mathbb{R}^{n}$ and $J(\bm{x})>$-$\infty$ for every $\bm{x}\in\mathbb{R}^{n}$). Consider the minimization problem $(\bm{x},t)\mapsto S(\bm{x},t)$ defined in Eq. 1 and its proximal operator $(\bm{x},t)\mapsto\text{prox}_{tJ}(\bm{x})$ defined in Eq. 2. We say a proper function $f_{t}\colon\mathbb{R}^{n}\to\mathbb{R}$ is a proximal operator of $tJ$ if $f_{t}(\bm{x})\in\text{prox}_{tJ}(\bm{x})$ for every $\bm{x}\in\mathbb{R}^{n}$. Gribonval and Nikolova [gribonval2020characterization] proved that proximal operators are characterized in terms of the function $\psi\colon\mathbb{R}^{n}\times[0,+\infty)\to\mathbb{R}\cup\{+\infty\}$ defined by
(6) $\psi(\bm{x},t)=\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t).$
Theorem 2.1.
A proper function $f_{t}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a proximal operator of $tJ$ if and only if $\bm{x}\mapsto\psi(\bm{x},t)$ is proper, lower semicontinuous, and convex and $f_{t}(\bm{x})\in\partial_{\bm{x}}\psi(\bm{x},t)$. Moreover, $f_{t}$ is uniformly Lipschitz continuous with constant $L>0$ if and only if $\bm{x}\mapsto(1-1/L)\left\|{\bm{x}}\right\|_{2}^{2}/2+tJ(\bm{x})$ is proper, lower semicontinuous and convex.
Proof 2.2.
See [gribonval2020characterization, Theorem 3 and Proposition 2]
The characterization of proximal operators in Theorem 2.1 is closely related to the concepts of semiconcave and semiconvex functions.
Definition 2.3.
Let $\mathcal{C}\subset\mathbb{R}^{n}$. We say $g\colon\mathcal{C}\to\mathbb{R}$ is $C$-semiconcave with $C\geqslant 0$ if it is continuous and
$\lambda g(\bm{x}_{1})+(1-\lambda)g(\bm{x}_{2})-g(\lambda\bm{x}_{1}+(1-\lambda)\bm{x}_{2})\leqslant\lambda(1-\lambda)C\left\|{\bm{x}_{1}-\bm{x}_{2}}\right\|_{2}^{2}$
for every $\bm{x}_{1},\bm{x}_{2}\in\mathcal{C}$ such that $\lambda\bm{x}_{1}+(1-\lambda)\bm{x}_{2}\subset\mathcal{C}$ and $\lambda\in[0,1]$. We say $g$ is semiconvex if g-g is semiconcave.
Remark 2.4.
It can be shown [cannarsa2004semiconcave, Chapter 1] that a function $g$ is $C$-semiconcave with $C\geqslant 0$ if and only if $\bm{x}\mapsto g(\bm{x})-\frac{C}{2}\left\|{\bm{x}}\right\|_{2}^{2}$ is concave, if and only if $g=g_{1}+g_{2}$, where $g_{1}$ is concave and $g_{2}\in C^{2}(\mathbb{R}^{n})$ with $\left\|{\nabla_{\bm{x}}^{2}g_{2}}\right\|_{\infty}\leqslant C$.
Combining formula Eq. 6Definition 2.3 and Remark 2.4, we find $\bm{x}\mapsto\psi(\bm{x},t)$ is convex if and only if $\bm{x}\mapsto tS(\bm{x},t)$ is semiconcave. We will see later that semiconcavity is an important concept in the theory of HJ PDEs for characterizing their generalized solutions. But before moving on to present some background on HJ PDEs, we give below an instructive example.
Example 2.5 (The negative absolute value prior).
Let J(x)=|x|J(x)=-|x| and consider the one-dimensional problem
$S(x,t)=\min_{y\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}-|y|\right\}.$
A global minimum $y^{*}$ of this problem satisfies the first-order optimality condition
$0\in(y^{*}-x)/t-\partial|y^{*}|\iff y^{*}\in\begin{cases}x+t,&\ \text{if $y^{*}>0$,}\\ [x-t,x+t]&\ \text{if $y^{*}=0$},\\ x-t,&\ \text{if $y^{*}<0$.}\end{cases}$
A short calculation shows that
[…]
Thus, a selection $f_{t}(x)\in\text{prox}_{tJ}(x)$ differs only at $x=0$. In any case, the function $x\mapsto\psi(x,t)$ in Theorem 2.1 and its subdifferential $x\mapsto\partial_{x}\psi(x,t)$ are given by
$\psi(x,t)=\frac{1}{2}x^{2}-tS(x,t)=\frac{1}{2}x^{2}+t|x|+\frac{t^{2}}{2}\quad\text{and}\quad\partial_{x}\psi(x)=\begin{cases}x+t,\,&\text{if $x>0$},\\ [-t,t],\,&\text{if $x=0$},\\ x-t,\,&\text{if $x<0$}.\end{cases}$
We see that any selection $f_{t}(x)\in\text{prox}_{tJ}(x)$ satisfies $f(x)\in\partial\psi(x,t)$.
2.2 Hamilton–Jacobi Equations
In this section, we briefly review some elements of the theory of HJ PDEs, including the method of characteristics, viscosity solutions of HJ PDEs, and the Lax–Oleinik formula, and discuss how these concepts tie together to proximal operators. The discussion is not comprehensive; see [evans2022partial] and references therein for a more detailed treatment. To ease the presentation, we consider only the first-order HJ PDEs Eq. 3.
2.2.1 Characteristic equations
The characteristic equations of Eq. 3 are given by the dynamical system
(8) $\begin{cases}\dot{\bm{x}}(t)&=\bm{p}(t),\\ \dot{\bm{p}}(t)&=0,\\ \dot{\bm{z}}(t)&=\frac{1}{2}\left\|{\bm{p}(t)}\right\|_{2}^{2},\end{cases}$
where $\bm{z}(t)=S(\bm{x}(t),t)$ and $\bm{x}(0)=J(\bm{x}(0))$. Here, $t\mapsto\bm{p}(t)$ is constant with $\bm{p}(t)\equiv\bm{p}(0)\in\mathbb{R}^{n}$. The characteristic line that arises from $\bm{x}(0)\in\mathbb{R}^{n}$ is $\bm{x}(t)=\bm{x}(0)+t\bm{p}(0)$, and so $\bm{z}(t)=\bm{z}(0)-\frac{1}{2}\left\|{\bm{p}(0)}\right\|_{2}^{2}$. Taken together, we find
$S(\bm{x}(t),t)=J(\bm{x}(0))+\frac{1}{2}\left\|{\bm{p}(0)}\right\|_{2}^{2}.$
Writing $\bm{x}(t)\equiv\bm{x}$ and $\bm{p}(0)=\nabla_{\bm{x}}S(\bm{x},t)$ (assuming formally that the spatial gradient exists at $\bm{x}$) then $\bm{x}(0)=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$, and so we find the representation
(9) $S(\bm{x},t)=\frac{1}{2t}\left\|{\nabla_{\bm{x}}S(\bm{x},t)}\right\|_{2}^{2}+J(\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)).$
This gives an implicit representation between $S$, its spatial gradient, and the initial data $J$. Next, we turn to the explicit representation of solutions to Eq. 3.
2.2.2 Viscosity solutions and the Lax–Oleinik formula
The initial value problem Eq. 3 (and HJ PDEs with general Hamiltonians) may not have a unique generalized solution, i.e., those satisfying the HJ PDE almost everywhere along with the initial condition $S(\bm{x},0)=J(\bm{x})$.
Example 2.6.
Let $J\equiv 0$ in Eq. 3 and take $n=1$. The corresponding HJ PDE has infinitely many solutions: For instance, the functions $S_{1}$ and $S_{2}$ given by
$S_{1}(x,t)=0,\quad S_{2}(x,t)=\begin{cases}0,\,&\text{if $|x|\geqslant t$},\\ x-t,\,&\text{if $0\leqslant x\leqslant t$},\\ -x-t,\,&\text{if $-t\leqslant x\leqslant 0$},\end{cases}$
satisfy $S_{1}(x,0)=S_{2}(x,0)=0$ and solve the corresponding HJ PDE.
The notion of viscosity solution was introduced in [crandall1983viscosity] to solve this problem. Under appropriate conditions (see [bardi1998hopf, crandall1992user, crandall1983viscosity]), the viscosity solution is unique and admits a representation formula. Specifically, for the initial value problem Eq. 3 with uniformly Lipschitz continuous initial data $J$, the unique viscosity solution is given by the Lax–Oleinik formula (with quadratic Hamiltonian)
(10) $S(\bm{x},t)=\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J(\bm{y})\right\}.$
The unique viscosity solution has two important properties. First, the function $\bm{x}\mapsto S(\bm{x},t)$ is (1/t)-semiconcave, that is, the function $\bm{x}\mapsto\psi(\bm{x},t)$ defined in Eq. 6 is convex, as per Theorem 2.1. Second, at any point of differentiability of $\bm{x}\mapsto S(\bm{x},t)$,
(11) $\nabla_{\bm{x}}S(\bm{x},t)=\frac{\bm{x}-f_{t}(\bm{x})}{t}\iff f_{t}(\bm{x})=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t),$
where $f_{t}(\bm{x})$ denote a global minimum in Eq. 10. Note that substituting this expression in formula Eq. 9 obtained from the characteristic equations yields Eq. 10, as expected.
Example 2.7 (The negative absolute value prior, continued.).
Let J(x)=|x|J(x)=-|x| in the HJ PDE Eq. 3. The function $J$ is uniformly Lipschitz continuous and, as such, the Lax–Oleinik formula $S(x,t)=-\frac{t}{2}-|x|$ is the unique viscosity solution of the corresponding HJ PDE. Note $\bm{x}\mapsto S(\bm{x},t)$ is differentiable everywhere except at $\bm{x}=0$ and $\text{prox}_{tJ}(\bm{x})=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$ everywhere except at $\bm{x}=0$ (see (7)).
In summary, a proper function $f_{t}$ is a proximal operator of $tJ$ whenever the function $(\bm{x},t)\mapsto S(\bm{x},t)$ is the viscosity solution of the HJ initial value problem Eq. 3. The minimization problem underlying $\text{prox}_{tJ}(\bm{x})$ is exactly the Lax–Oleinik representation formula of the viscosity solution of Eq. 3. We will see in the next section how to leverage these connections for learning the prior when $\bm{x}\mapsto J(\bm{x})$ is not available but $(\bm{x},t)\mapsto S(\bm{x},t)$ is available. But before proceeding, we briefly review convex neural networks, which will be used later in this work.
2.3 Convex neural networks
Convex Neural Networks, specifically Input Convex Neural Networks (ICNN), were introduced by [amos2017input] to allow for the efficient optimization of neural networks within structured prediction and reinforcement learning tasks. The core premise of an ICNN is to constrain the network architecture such that the output is a convex function with respect to the input.
To achieve convexity, the network typically employs a recursive structure for $k=0,\dots,j-1$
(12) $\bm{z}_{k+1}=g(\bm{W}_{k}\bm{z}_{k}+\bm{H}_{k}y+\bm{b}_{k}),f(\bm{y};\theta)=\bm{z}_{j},$
where $\bm{y}$, $\bm{z}_{k}$ represent the input to the network and the hidden features at layer $k$, respectively, and $g$ is the activation function. To guarantee the convexity of the output with respect to the input $\bm{y}$, specific constraints are imposed on the parameters and the activation function, which are (i) the weights $\bm{W}_{k}$, which connect the previous hidden layer to the current one, must be non-negative ($\bm{W}_{k}\geqslant 0$), and (ii) the activation function $g$ must be convex and non-decreasing [fang2024whats].
Following [fang2024whats, Proposition  3.1], Fang et al. leverage the ICNN architecture and the characterization of proximal operators to develop Learned Proximal Networks (LPN) for inverse problems. LPNs require stricter conditions than standard ICNNs. While standard ICNNs often use ReLU activation, LPNs require the activation function $g$ to be twice continuously differentiable. This smoothness is essential to ensure that the proximal operator is the gradient of a twice continuous differentiable function [gribonval2020characterization, Theorem 2]. Consequently, LPNs typically utilize smooth activations like the softplus function, a $\beta$-smooth approximation of ReLU [fang2024whats, Section 3].
3 Connections between learning priors and the inverse problem for Hamilton–Jacobi equations
In this section, we discuss the inverse problem of learning the prior in the proximal operator Eq. 2: given $t>0$ and some function $\bm{x}\mapsto S(\bm{x},t)$, assess whether there exists a a prior function $J$ that can recover $\bm{x}\mapsto S(\bm{x},t)$ and, if so, estimate it. Due to the connections between proximal operators and HJ equations, as discussed in Subsections 2.12.2, our starting point will be to discuss the inverse problem from the point of view of HJ equations.
We summarize in the next subsection some of the main results for this problem, based on the results of [esteve2020inverse] and other related works [claudel2011convex, colombo2020initial, misztela2020initial].
3.1 Reachability and inverse problems for Hamilton–Jacobi equations
We consider the inverse problem associated to the HJ initial value problem Eq. 3: given $t>0$ and a function $(\bm{x},t)\mapsto S(\bm{x},t)$, find the set of initial data $J\colon\mathbb{R}^{n}\ \to\mathbb{R}$ such that the viscosity solution of Eq. 3 is $S(\bm{x},t)$. That is, we wish to find the set
(13) $I_{t}(S)$ $\coloneqq\{\text{$J\colon\mathbb{R}^{n}\to\mathbb{R}$ is uniformly Lipschitz continuous}$
[…]
We say the function $(\bm{x},t)\mapsto S(\bm{x},t)$ is reachable if the set $I_{t}(S)$ is nonempty. The main reachability result for the initial value problem Eq. 3 is the following:
Theorem 3.1.
Suppose $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous. Then the set $I_{t}(S)$ defined in Eq. 13 is nonempty if and only if $\bm{x}\mapsto tS(\bm{x},t)$ is semiconcave.
Proof 3.2.
This follows from [esteve2020inverse, Theorem 2.2, Theorem 6.1, and Definition 6.2].
Now, assume $(\bm{x},t)\mapsto S(\bm{x},t)$ is reachable. What can we say about $I_{t}(S)$? Since we evolve forward in time $J$ from 0 to $t$ via Eq. 3 to obtain $S(\bm{x},t)$, a natural approach then is to do the opposite: evolve backward in time $S(\bm{x},t)$ from $t$ to 0 via an appropriate HJ PDE. That is, we consider the terminal value problem
(14) $\begin{dcases}\frac{\partial\bm{w}}{\partial\tau}(\bm{y},\tau)+\frac{1}{2}\left\|{\nabla_{\bm{y}}\bm{w}(\bm{y},\tau)}\right\|_{2}^{2}=0&\ (\bm{y},\tau)\in\mathbb{R}^{n}\times[0,t),\\ \bm{w}(\bm{y},t)=S(\bm{y},t),&\ \bm{y}\in\mathbb{R}^{n}.\end{dcases}$
Under certain conditions, the HJ PDE Eq. 14 has a unique viscosity solution:
Theorem 3.3.
Suppose $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and semiconcave. Then the HJ PDE Eq. 14 has a unique viscosity solution given by
(15) $\bm{w}(\bm{y},\tau)=\sup_{\bm{x}\in\mathbb{R}^{n}}\left\{S(\bm{x},t)-\frac{1}{2\tau}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}\right\}.$
Moreover, the function $\bm{y}\mapsto\tau\bm{w}(\bm{y},\tau)$ is semiconvex with unit constant.
Proof 3.4.
See [barron1999regularity, Section 4, Equation 4.4.2] and [cannarsa2004semiconcave, Chapter 1].
The viscosity solution of (14) is sometimes called the backward viscosity solution (BVS) to distinguish it from its forward counterpart given by Eq. 3. The BVS at $\tau=0$ corresponds to fully evolving backward in time the function $\bm{x}\mapsto S(\bm{x},t)$. In what follows, we write $J_{\text{BVS}}\coloneqq\bm{w}(\cdot,0)$. We can use Eq. 6 to write
(16) $tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}=\sup_{\bm{x}\in\mathbb{R}^{n}}\left\{\langle\bm{x},\bm{y}\rangle-\psi(\bm{x},t)\right\}.$
The right hand side is the convex conjugate of $\bm{x}\mapsto\psi(\bm{x},t)$ evaluated at $\bm{x}$, which is well-defined because $\bm{x}\mapsto\psi(\bm{x},t)$ is proper, lower semicontinuous and convex.
Theorem 3.3 suggests that $J_{\text{BVS}}$ is an initial condition that can reach $\bm{x}\mapsto S(\bm{x},t)$. The next result stipulates that this is correct and that it is “optimal”, in the sense that it bounds from below for any other reachable initial condition $J\in I_{t}(S)$.
Theorem 3.5.
Let $J_{\text{BVS}}$ denote the solution of the backward HJ terminal value problem 14 at time $\tau=0$. Then $J\in I_{t}(S)$ if and only if
[…]
$X_{t}(S)\coloneqq\left\{\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t):\text{$\bm{x}\mapsto S(\bm{x},t)$ is differentiable at $\bm{x}\in\mathbb{R}^{n}$}\right\}.$
Proof 3.6.
See [esteve2020inverse, Theorems 2.3 and 2.4].
Theorem 3.5 is a fundamental consequence of the semiconcavity of $\bm{x}\mapsto S(\bm{x},t)$, which regularizes the backward viscosity solution of Eq. 14. We illustrate this below with the negative absolute value prior.
Example 3.7 (The negative absolute value prior, continued.).
Let J(x)=|x|J(x)=-|x| in the HJ PDE Eq. 3. Recall that its unique viscosity solution is given by the Lax–Oleinik formula $S(x,t)=-\frac{t}{2}-|x|$. We now compute the corresponding unique backward viscosity solution to the terminal-value problem Eq. 14. The solution is well-defined because $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and concave. We have
$J_{\text{BVS}}(x)=\sup_{\bm{y}\in\mathbb{R}}\left\{-\frac{t}{2}-|y|-\frac{1}{2t}(x-y)^{2}\right\}=-\frac{t}{2}-\inf_{\bm{y}\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}+|y|\right\}.$
The infimum on the right hand side corresponds to the proximal operator of the function $\bm{y}\mapsto|y|$, which is the soft-thresholding operator:
(17) $\operatorname*{arg\,min}_{y\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}+|y|\right\}=\begin{cases}x-t,\,&\text{if $x>t$},\\ 0,\,&\text{if $x\in[-t,t]$},\\ x+t,\,&\text{if $x<-t$}.\end{cases}$
This gives
$J_{\text{BVS}}(x)=\begin{cases}-|x|,\,&\text{if $x<-t$ or $x>t$},\\ -\frac{t}{2}-\frac{x^{2}}{2t},\,&\text{if $x\in[-t,t]$}.\end{cases}$
Here, a simple calculation shows $\mathcal{X}_{t}(S)=(-\infty,-t]\cup[t,+\infty)$, and we find $J(x)>J_{\text{BVS}}(x)$ on (t,t)(-t,t), as expected from Theorem 3.5. Moreover,
(18) $tJ_{\text{BVS}}(x)+\frac{1}{2}x^{2}=\begin{cases}\frac{1}{2}(x-t)^{2}-\frac{t^{2}}{2},\,&\text{if $x>t$},\\ -\frac{t^{2}}{2},\,&\text{if $x\in[-t,t]$},\\ \frac{1}{2}(x+t)^{2}-\frac{t^{2}}{2},\,&\text{if $x<-t$},\end{cases}$
and we observe $x\mapsto tJ_{\text{BVS}}(x)+\frac{1}{2}x^{2}$ is convex, as expected from Theorem 3.3.
The results here apply when the function $\bm{x}\mapsto S(\bm{x},t)$ is known. Can anything be said when only a finite set of values of $\bm{x}\mapsto S(\bm{x},t)$ is available?
4 Learning priors and the inverse problem for Hamilton–Jacobi Equations with incomplete information
In this section, we consider the inverse problem associated to the HJ initial value problem Eq. 3 with incomplete data: given $t>0$ and samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$, learn the prior $J$ that “best” recovers the observations $\{S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$. Assuming $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous, recall from Theorem 3.1 that $\bm{x}\mapsto S(\bm{x},t)$ is reachable if and only if it is semiconcave. If reachable, the backward viscosity solution $\bm{x}\mapsto J_{\text{BVS}}(\bm{x})$ of the HJ terminal value problem Eq. 14 then recovers $(\bm{x},t)\mapsto S(\bm{x},t)$. Thus, we will focus on studying how to learn $J_{\text{BVS}}$.
As a first step, note that the unique minimum in the Lax–Oleinik formula Eq. 10 at the sample $\bm{x}=\bm{x}_{k}$ can be represented via Eq. 11:
(19) $S(\bm{x}_{k},t)=\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}+J(\bm{y}_{k}),\ \text{where}\ \bm{y}_{k}=\bm{x}_{k}-t\nabla_{\bm{x}}S(\bm{x}_{k},t).$
Moreover, Theorem 3.5 implies $J(\bm{y})\geqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, and in particular we have $J(\bm{y}_{k})=J_{\text{BVS}}(\bm{y}_{k})$. Taken together, this gives samples $\{\bm{y}_{k},J(\bm{y}_{k})\}_{k=1}^{K}$ of the true underlying prior $J$ (and $J_{\text{BVS}}$).
As a second step, recall the function $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is convex by Theorem Theorem 3.3. Thus instead of learning $J_{\text{BVS}}$, we can equivalently learn the convex function $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ using the samples $\{\bm{y}_{k},J(\bm{y}_{k})\}_{k=1}^{K}$. Before we propose methods for learning this convex function numerically using deep learning and convex neural networks (the topic of Section 5), we discuss below different approaches for learning $J_{\text{BVS}}$, in particular by approximating $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ from below as a piecewise convex function at the points $\left\{\bm{y}_{k}\right\}_{k=1}^{K}$. We consider this below in Section 4.1. This approximation problem, it turns out, is related to max-plus algebra theory for approximating solutions to HJ PDEs [akian2006max, fleming2000max, gaubert2011curse]; we discuss this in Section 4.2. We specialize these results to the problem of approximating $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ in general in Section 4.3.
4.1 Piecewise approximations
We consider here piecewise approximations of the prior $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})$ using the samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ and formula Eq. 19. We consider first using a piecewise affine minorant (PAM) approximation, and then, assuming some regularity on $J_{\text{BVS}}$, using a piecewise quadratic minorant (PQM) approximation.
4.1.1 Piecewise affine approximation
We first consider the PAM approximation of the convex function $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$:
(20) $tJ_{\text{PAM}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}\coloneqq\max_{k\in\{1,\dots,K\}}\left\{tJ_{\text{BVS}}(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{y}_{k}}\right\|_{2}^{2}+\left\langle\bm{x}_{k},\bm{y}-\bm{y}_{k}\right\rangle\right\}.$
Then $J_{\text{PAM}}(\bm{y})\leqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, with $J_{\text{PAM}}(\bm{y}_{k})=J_{\text{BVS}}(\bm{y}_{k})$ at each $k\in\{1\,\dots,K\}$. A short calculation gives
(21) $tJ_{\text{PAM}}(\bm{y})=\max_{k\in\{1,\dots,K\}}\left\{tJ_{\text{BVS}}(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}\right\}.$
How good is $J_{\text{PAM}}$ as an initial condition for the HJ PDE Eq. 3? A short calculation yields
(22) $\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PAM}}(\bm{y})\right\}=\begin{cases}S(\bm{x}_{k},t)&\ \text{if $\bm{x}=\bm{x}_{k}$, $k\in\{1,\dots,K\}$},\\ +\infty,&\ \text{otherwise},\end{cases}$
The PAM approximation recovers $\{S(\bm{x}_{k},t)\}_{k=1}^{K}$ and nothing else. It is therefore generally inadequate to use as a meaningful learned prior or estimate of $J_{\text{BVS}}$.
4.1.2 Piecewise quadratic approximation
Assume now $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})$ is semiconvex with constant $(1-\alpha)$, $\alpha>0$. Then $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is $(1-\alpha)$ strongly convex, and we can approximate it via its piecewise quadratic minorant (PQM):
$tJ_{\text{PQM}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ $\coloneqq\max_{k\in\{1,\dots,K\}}\biggl\{tJ_{\text{BVS}}(\bm{y}_{k})\ +$
$\qquad\qquad\left.\frac{1}{2}\left\|{\bm{y}_{k}}\right\|_{2}^{2}+\langle\bm{x}_{k},\bm{y}-\bm{y}_{k}\rangle+\frac{\alpha}{2}\left\|{\bm{y}-\bm{y}_{k}}\right\|_{2}^{2}\right\}.$
Then, $J_{\text{PQM}}(\bm{y})\leqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, with $J_{\text{PQM}}(\bm{y})=J_{\text{BVS}}(\bm{y}_{k})$ at each $k\in\{1,\dots,K\}$. Moreover, a short calculation gives
$tJ_{\text{PQM}}(\bm{y})=\max_{k\in\{1,\dots,K\}}\left\{J(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}+\frac{\alpha}{2}\left\|{\bm{y}-\bm{y}_{k}}\right\|_{2}^{2}\right\}.$
How good is $J_{\text{PQM}}$ as an initial condition for the HJ PDE Eq. 3? A formal calculation, relegated to Appendix A, yields
(23) $\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PQM}}(\bm{y})\right\}=\frac{1}{2t}\left\|{\bm{x}-\bm{y}_{k}}\right\|_{2}^{2}+\frac{1}{2t\alpha}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}$
for some $k\in\{1,\dots,K\}$. The PQM approximation now leads to an approximation of $(\bm{x},t)\mapsto S(\bm{x},t)$ that recovers $\{S(\bm{x}_{k},t)\}$ and is finite everywhere. How accurate it is, however? In the next section, we describe how max-plus algebra theory [akian2006max, fleming2000max, gaubert2011curse] can be used to quantify the approximation errors more precisely.
4.2 Max-plus algebra theory for Hamilton–Jacobi PDEs and approximation results
We consider here max-plus algebra techniques for approximating solutions to certain HJ PDEs. Let $\alpha>0$ and let $\Psi\colon\mathbb{R}^{n}\to\mathbb{R}$ denote a $(1-\alpha)$-semiconvex function. Following [gaubert2011curse, Section III], we approximate $\Psi$ using $K$ vectors $\{\bm{p}_{k}\}_{k=1}^{K}\subset\mathbb{R}^{n}$ with $K$ semiconvex functions $\bm{y}\mapsto\langle\bm{p}_{k},\bm{y}\rangle-\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ and a function $a\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$:
(24) $\Psi_{\text{MP}}(\bm{y})\coloneqq\max_{k\in\{1,\dots,K\}}\left\{\langle\bm{p}_{k},\bm{y}\rangle-\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}-a(\bm{p}_{k})\right\}.$
Here, we suppose the vectors $\{\bm{p}_{k}\}_{k=1}^{K}$ and $\bm{p}\mapsto a(\bm{p})$ are selected so that $\Psi(\bm{y})\geqslant\Psi_{\text{MP}}(\bm{y})$. As discussed in Section 4.1, such a selection is possible via the PQM of the $(1-\alpha)$-strongly convex function $\bm{y}\mapsto\Psi(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$. Let $\mathcal{Y}$ denote a full dimensional compact, convex subset of $\mathbb{R}^{n}$ and consider the $L_{\infty}$ error
$\epsilon_{\infty}(\Psi,K,\mathcal{Y},\Psi_{\text{MP}})\coloneqq\sup_{\bm{y}\in\mathcal{Y}}|\Psi(\bm{y})-\Psi_{\text{MP}}(\bm{y})|.$
Furthermore, we define the corresponding minimal $L_{\infty}$ error as
$\delta_{\infty}(\Psi,K,\mathcal{Y})=\inf_{\Psi_{\text{MP}}\leqslant\Psi}\epsilon_{\infty}(\Psi,K,\mathcal{Y},\Psi_{\text{MP}}).$
The following result from [gaubert2011curse], stipulates that whatever vectors $\{\bm{p}_{k}\}_{k=1}^{K}$ and function $\bm{p}\mapsto a(\bm{p})$ are used to approximate $\Psi$, the minimal $L_{\infty}$ error scales as an inverse power law in $K$ and the dimension $n$ in the limit $K\to+\infty$.
Theorem 4.1 (Gaubert et al. (2011)).
Let $\alpha>0$, and let $\mathcal{Y}$ denote a full-dimensional compact, convex subset of $\mathbb{R}^{n}$. If $\Psi\colon\mathbb{R}^{n}\to\mathbb{R}$ is twice continuously differentiable and $(1-\alpha)$ semiconvex, then there is a constant $\beta(n)>0$ such that
(25) $\delta_{\infty}(\Psi,K,\mathcal{Y})\sim\beta(n)\left(\frac{1}{K}\int_{\mathcal{Y}}\left(\det\left(\nabla_{\bm{y}}^{2}\Psi(\bm{y})+\bm{I}_{n\times n}\right)\right)^{\frac{1}{2}}\mathop{}\!d\bm{y}\right)^{2/n}\ \text{as $K\to+\infty$.}$
Thus the minimal $L_{\infty}$ error is $\Omega(1/K^{2/n})$ as $K\to+\infty$, though the error is smaller the closer the Hessian matrix $\nabla_{\bm{y}}^{2}\Psi(\bm{y})$ is to the identity matrix $\bm{I}_{n\times n}$.
4.3 Applications to the inverse problem for Hamilton–Jacobi Equations
We now return to the setting of Section 4.1, namely to approximate the backward viscosity solution $\bm{x}\mapsto J_{\text{BVS}}(\bm{x})$ using samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ and assuming, additionally, that $J_{\text{BVS}}$ is sufficiently “regularized”. Max-plus algebra theory provides us with the following approximation result in the “large sample size” setting.
Theorem 4.2.
Let $t>0$ and assume $tJ_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$-semiconvex with $\alpha>0$. Let $\mathcal{Y}$ denote a full-dimensional compact, convex set of $\mathbb{R}^{n}$ and let $\tilde{J}\in I_{t}(S)$ denote a function that can reach $\bm{x}\mapsto S(\bm{x},t)$. Then there exists a constant $\beta(n)$ depending only on $n$ such that
(26) $\delta_{\infty}(tJ_{\text{BVS}},K,\mathcal{Y})\sim\beta(n)\left(\frac{1}{K}\int_{\mathcal{Y}}\left(\det\left(t\nabla_{\bm{y}}^{2}J_{\text{BVS}}(\bm{y})+\bm{I}_{n\times n}\right)\right)^{\frac{1}{2}}\mathop{}\!d\bm{y}\right)^{2/n}$
as $K\to+\infty$, and
(27) $\delta_{\infty}(J_{\text{BVS}},K,\mathcal{Y})\leqslant t\sup_{\bm{y}\in\mathcal{Y}}|\tilde{J}(\bm{y})-J_{\text{PQM}}(\bm{y})|.$
Proof 4.3.
The asymptotic limit Eq. 26 follows immediately from Theorem 4.1 because $J_{\text{BVS}}$ satisfies all its assumptions. For inequality Eq. 27, first note Theorem 3.5 implies $\tilde{J}(\bm{y})\geqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, with equality for every $\bm{y}\in\mathbb{R}^{n}$ for which $\bm{y}=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$ for some $\bm{x}\in\mathbb{R}^{n}$. Thus
$t\tilde{J}(\bm{y})-tJ_{\text{BVS}}(\bm{y})=(t\tilde{J}(\bm{y})-tJ_{\text{PQM}}(\bm{y}))+(tJ_{\text{PQM}}(\bm{y})-tJ_{\text{BVS}}(\bm{y}))\geqslant 0,$
which we rearrange to get
$tJ_{\text{BVS}}(\bm{y})-tJ_{\text{PQM}}(\bm{y})\leqslant t\tilde{J}(\bm{y})-tJ_{\text{PQM}}(\bm{y}).$
Since the set $\mathcal{Y}$ is a compact and convex set, $\sup_{\bm{y}\in\mathcal{Y}}|tJ_{\text{BVS}}(\bm{y})-tJ_{\text{PQM}}(\bm{y})|$ is finite and attained in $\mathcal{Y}$, say at $\bm{y}^{*}$. Combining this with the inequality above yields
$t\sup_{\bm{y}\in\mathcal{Y}}|J_{\text{BVS}}(\bm{y})-J_{\text{PQM}}(\bm{y})|\leqslant t\tilde{J}(\bm{y}^{*})-tJ_{\text{PQM}}(\bm{y}^{*})\leqslant t\sup_{\bm{y}\in\mathcal{Y}}|\tilde{J}(\bm{y})-J_{\text{PQM}}(\bm{y})|.$
Finally, since $J_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$ semiconvex with $\alpha>0$, we can invoke Theorem 4.1 with $\Psi\equiv J_{\text{BVS}}$ to get inequality Eq. 27. This concludes the proof.
Thus $J_{\text{PQM}}$ approximates $J_{\text{BVS}}$ from below in $\Omega(1/K^{n/2})$ as $K\to+\infty$, and the $L_{\infty}$ error of any reachable function $\tilde{J}\in I_{t}(S)$ with respect to $J_{\text{PQM}}$ is bounded from below by this minimal error $\delta_{\infty}(tJ_{\text{BVS}},K,\mathcal{Y})$. Thus if there are “enough” samples approximating “well” $\bm{x}\mapsto S(\bm{x},t)$, then the PQM approximation will converge. In addition, its error rate may be small if the determinant term in Eq. 26 is close to zero.
What can be said when $tJ_{\text{BVS}}$ is not $(1-\alpha)$-semiconvex? Without further information on the prior $J$ or other regularizing properties of $tJ_{\text{BVS}}$, the answer is unclear, at least to the best of the knowledge of the authors. Deep learning methods can provide a way to train a meaningful prior to approximate $tJ_{\text{BVS}}$, e.g., by fitting well or perfectly a convex neural network $\bm{y}\mapsto G_{\theta}(\bm{y})$ to the function $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ and enforcing enough regularity [^2] It is enough, e.g., to enforce $\bm{y}\mapsto G_{\theta}(\bm{y})/t-\frac{1}{2t}\left\|{\bm{y}}\right\|_{2}^{2}$ to be uniformly Lipschitz continuous, since then the infimum problem would correspond to the Lax–Oleinik formula to the solution of an appropriate forward HJ PDE [bardi1998hopf, crandall1992user, crandall1983viscosity]. to ensure
$\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+\frac{1}{t}\left(G_{\theta}(\bm{y})-\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}\right)\right\}$
is finite for every $\bm{x}\in\mathbb{R}^{n}$. In the next section, we present some numerical results to test this proposed deep neural network approach.
5 Numerical results
In this section, we leverage the theory of viscosity solutions of HJ PDEs to learn proximal operators of certain (non)convex priors. We used Learned Proximal Networks (LPNs) [fang2024whats] for this, which are designed to learn the proximal operator as the gradient of a convex potential. These experiments investigate their behavior when trained on data generated from fundamentally nonconvex and concave landscapes. All experiments used the official LPN implementation [fang2024whats].
We used LPNs with $2$ layers and $256$ hidden units using Softplus activation ($\beta=5$) to ensure $C^{2}$ smoothness. The models were trained via the Adam optimizer with a starting learning rate of $10^{-3}$ and decreased by a factor of $10^{-1}$ at every $10^{5}$ epochs for a total of $5\times 10^{5}$ epochs. We generated the data for all experiments as follow: $N$ samples $\{\bm{x}_{k}\}_{k=1}^{N}$ were drawn uniformly from the hypercube $a,a]^{d}$, where $a=4$ and $d=2,4,8,16,32$ and $64$. We chose $N=3\times 10^{4}$ $d=2$ and $4$, $N=3\times 10^{4}$ for $d=8$ and $16$, and $N=4\times 10^{4}$ for $d=32$ and $64$. In each experiment, we evaluated the minimum value of the proximal operator $S(\bm{x}_{k},1)$ at every sample.
The first phase of the experiment focused on learning the value function $S(\bm{x},1)$ associated with the HJ equation Eq. 3 with initial data $J$ at $t=1$. We encoded it into a convex potential function $\psi_{\text{NN}}(\bm{y};\theta)$, which we trained by minimizing the MSE loss
$\min_{\theta}\frac{1}{K}\sum_{k=1}^{K}\left(\psi_{\text{NN}}(\bm{x}_{k};\theta)-\left(\frac{1}{2}\left\|{\bm{x}_{k}}\right\|_{2}^{2}-S(\bm{x}_{k},1)\right)\right)^{2}.$
After training, we evaluated the trained value function using $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-\psi_{\text{NN}}(\bm{x};\theta)$.
[…]
$\bm{y}_{k}=\nabla_{\bm{x}}\psi_{\text{NN}}(\bm{x}_{k};\theta),\quad J_{\text{BVS}}(\bm{y}_{k})\approx G_{k}\coloneqq\langle\bm{x}_{x},\bm{y}_{k}\rangle-\psi_{\text{NN}}(\bm{x}_{k},1)-\frac{1}{2}\left\|{\bm{y}_{k}}\right\|_{2}^{2}.$
We then trained the second LPN by minimizing the MSE loss
$\min_{\tilde{\theta}}\frac{1}{K}\sum_{k=1}^{K}\left(J_{\text{NN}}(\bm{y}_{k};\tilde{\theta})-G_{k}\right)^{2}.$
[…]
5.1 Convex prior: the $\ell_{1}$ norm
We begin with the convex prior $J(\bm{x})=\left\|{\bm{x}}\right\|_{1}$. For this example, the value function $S(\bm{x},t)$ can be evaluated using the soft-thresholding operator Eq. 17. The errors with respect the exact solutions of the first and second LPN, are presented in the Table 1. Some plots of cross sections of $\psi_{\text{NN}}$ with respect to the true convex function $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-S(\bm{x},1)$ are presented in Fig. 1 below. The results demonstrate that the first LPN learns fairly well the forward solution $\bm{x}\mapsto S(\bm{x},t)$ from samples of $\bm{x}$ and $S(\bm{x},1)$ even in high dimensions, though the MSE taper off as the dimension $d$ increases.
We wish to note that in dimension 64, the MSE for $\psi_{\text{NN}}$ reduced significantly compared to the previous dimensions. We are unsure as to why this is the case. It is possible that using a deeper neural network or incorporating more nodes may improve performance. The second LPN learns the true prior $\bm{x}\mapsto\left\|{\bm{x}}\right\|_{1}$ also fairly well, especially since it learning from the first, approximate LPN $\psi_{\text{NN}}$, without recourse to samples of the gradient $\bm{x}\mapsto\nabla_{\bm{x}}S(\bm{x},t)$.
Table 1: Mean square errors of LPN $\psi_{\text{NN}}$ and prior $J_{\text{NN}}$ with 2 layers and 256 neurons in the convex L1 prior example.
Dimension LPN ($\psi_{\text{NN}}$) Prior ($J_{\text{NN}}$)
Mean Square Errors 2D 1.04E51.04E-5 3.33E53.33E-5
4D 2.97E52.97E-5 2.17E42.17E-4
8D 1.05E41.05E-4 7.25E47.25E-4
16D 5.27E35.27E-3 2.11E32.11E-3
32D 1.6E11.6E-1 4.03E24.03E-2
64D 2.89E62.89E-6 2.69E32.69E-3
Figure 1: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
5.2 Non-convex priors
A min-plus algebra prior
For this example, we use a min-plus prior:
$J(\bm{x})=\min\left(\frac{1}{2\sigma_{1}}\left\|{\bm{x}-\mu_{1}}\right\|_{2}^{2},\frac{1}{2\sigma_{2}}\left\|{\bm{x}-\mu_{2}}\right\|_{2}^{2}\right).$
We used $\mu_{1}=(1,0,\dots,0)$, $\mu_{2}=\bm{1}/\sqrt{n}$, and $\sigma_{1}=\sigma_{2}=1.0$. We chose this prior despite it not being uniformly Lipschitz continuous, an assumption needed for having a unique viscosity solution to the HJ PDE, to check if an LPN could learn the prior regardless. We obtained the following results, presented in Table 2 and Fig. 2 and  3. The results are comparable to those obtained for the $\ell_{1}$ prior in Section 5.1. The MSE for the $d=64$ case also decreases dramatically, as in the $\ell_{1}$-convex prior, and we are unsure why this is the case. In any case, the second LPN performed reasonably well up in high dimensions.
Table 2: Mean square errors of LPN $\psi$ and prior $J_{\text{NN}}$ with 2 layers and 256 neurons in the min-plus example.
Dimension LPN ($\psi_{\text{NN}}$) Prior ($J_{\text{NN}}$)
Mean Square Errors 2D 3.33E63.33E-6 5.73E75.73E-7
4D 7.64E67.64E-6 4.92E64.92E-6
8D 3.64E53.64E-5 1.20E41.20E-4
16D 1.99E41.99E-4 3.44E43.44E-4
32D 1.16E31.16E-3 1.33E31.33E-3
64D 2.32E92.32E-9 5.21E55.21E-5
Figure 2: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
Figure 3: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension $32$ (left). The cross sections of the prior function learned from our trained second LPN method (right).
5.3 Concave prior
For this example, we use the concave prior
$J(\bm{x})=-\left\|{\bm{x}}\right\|_{2}^{2}/4.$
This is a challenging example because, as the min-plus prior, it is not uniformly Lipschitz continuous, although numerically since we constrain the domain to $4,4]^{d}$, we can get around this by “Huberizing” the prior outside $4,4]^{d}$. For this prior, we have $S(\bm{x},t)=-\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}$. After training, we obtained the results presented in Table 3 and Figs. 4. The results are similar to those obtained for the two previous prior, though we found this prior slightly more challenging to train. Again, for dimension $d=64$, the MSE for the first LPN decreased by a large factor.
Table 3: Mean square errors of LPN $\psi_{\text{NN}}$ and prior $J_{\text{NN}}$ with 2 layers and 256 neurons in the concave prior example.
Dimension LPN ($\psi_{\text{NN}}$) Prior ($J_{\text{NN}}$)
Mean Square Errors 2D 7.00E77.00E-7 1.57E61.57E-6
4D 2.74E52.74E-5 7.70E57.70E-5
8D 5.58E45.58E-4 7.91E47.91E-4
16D 3.69E33.69E-3 3.28E33.28E-3
32D 8.70E28.70E-2 3.01E23.01E-2
64D 6.23E66.23E-6 1.87E31.87E-3
Figure 4: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
5.4 Negative $\ell_{1}$ norm
We finish with the nonconvex prior $J(\bm{x})=-\left\|{\bm{x}}\right\|_{1}$. As per examples 2.5 with $n=1$, we have
$S(\bm{x},t)=-\frac{n}{2}-\left\|{\bm{x}}\right\|_{1}.$
We present the results after training in Table 4 and some plots in Fig. 5. This example turned out to be very challenging; it did well up to and including dimension 8, but did not perform very well in higher dimensions. The MSE for the first LPN decreased by a large factor in $d=64$ compared to $d=32$, and again we are unsure why this is the case. Nonetheless, Fig. 5 shows that the learned prior turned out to approximate the true backward viscosity solution given by Eq. 18, as expected from the theory of backward viscosity solutions.
In addition to our numerical experiments, we show below the performance of the regularized inversion method proposed in [fang2024whats] to estimate $J$ from the first LPN [^3] Find $\bm{y}$ such that $\text{prox}_{J}(\bm{y})=\bm{x}$ by minimizing $\psi(\bm{y})+\frac{\alpha}{2}\left\|{\bm{y}}\right\|_{2}^{2}-\langle\bm{x},\bm{y}\rangle)$ w.r.t. $\bm{y}$ with $\alpha\geqslant 0$.. As Fig. 7 shows, the regularized inversion method performed very poorly and we thus do not comment on it further.
Table 4: Mean square errors of LPN $\psi$ and prior $J$ with 2 layers and 256 neurons in the negative $L_{1}$ norm examples.
Dimension LPN ($\psi$) Prior ($J$)
Mean Square Errors 2D 6.59E56.59E-5 5.20E65.20E-6
4D 3.15E43.15E-4 3.17E53.17E-5
8D 2.12E32.12E-3 2.94E42.94E-4
16D 8.01E38.01E-3 4.49E24.49E-2
32D 1.55E11.55E-1 2.29E22.29E-2
64D 6.42E46.42E-4 4.49E34.49E-3
Figure 5: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 44 (left). The cross sections of the prior function learned from our trained second LPN method (right).
Figure 6: The cross sections of the convex function $\psi_{\text{NN}}(\bm{x};\theta)$ for dimension 88 (left). The cross sections of the prior function learned from our trained second LPN method (right).
Figure 7: The cross sections of the prior function learned from the regularized inversion method proposed in [fang2024whats] Dimension 44 (left) and dimension 88 (right).
6 Discussion
In this work, we leveraged the theory of viscosity solutions of HJ PDEs and recent work on proximal operators to develop novel deep learning numerical methods for learning the underlying prior of the proximal operator Eq. 2 from data. Our approach built on existing connections between proximal operators and HJ PDEs, namely that proximal operators are obtained from solutions to a class of HJ PDEs, and on the theory for the inverse problem for HJ equations. As discussed in Section 3, while there may be infinitely many priors that can recover Eq. 1, there is a natural choice, obtained by reversing the time in the HJ PDE Eq. 14 and using the value of the proximal operator $(\bm{x},t)\mapsto S(\bm{x},t)$ as initial condition. The resulting backward viscosity solution yields a semiconvex prior $J_{\text{BVS}}$ that can reconstruct the value function $(\bm{x},t)\mapsto S(\bm{x},t)$ underlying proximal operator. We considered the special case where only samples of the proximal operators and its values were available in Section Section 4, and used techniques from max-plus algebra to derive some characterizations and errors property of $J_{\text{BVS}}$ with respect to convex functions approximating it from above. Finally, in Section 5 we presented numerical experiments to learn both the forward solution $\bm{x}\mapsto S(\bm{x},t)$ and the prior $J_{\text{BVS}}$ by training a convex neural network, specifically a learned proximal network from data $\{\bm{x}_{k},S(\bm{x}_{k},t)\}_{k=1}^{K}$. We presented several numerical results that demonstrate the efficiency of our proposed method in high dimensions.
While this work focused on proximal operators, we expect our approach can be extended to a broad class of Bregman divergences, as recent results in the theory of inverse problems for HJ equations suggest [esteve2020inverse]. Another potential direction would be in the case where the value of the proximal operator $(\bm{x},t)\mapsto S(\bm{x},t)$ is known to learn the prior $J$ using Monte Carlo sampling strategies, as recently proposed in [park2025neural] for the forward problem of HJ equations (i.e., learning $(\bm{x},t)\mapsto S(\bm{x},t)$ from known $J$). In the longer term, it would be interesting to devise similar deep learning methods for the inverse problem of HJ equations with possibly time- or state-dependent Hamiltonians, relevant to optimal control problems.
Appendix A Formal calculation of Eq. 23
Formally, we have
$t\nabla_{\bm{y}}J_{\text{PQM}}(\bm{y})=\alpha(\bm{y}-\bm{y}_{k})+\bm{x}_{k}-\bm{y}$
for some $k\in\{1,\dots,K\}$. Similarly,
$\hat{\bm{y}}=\operatorname*{arg\,min}_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PQM}}(\bm{y})\right\}$ $\iff\bm{0}=\frac{\hat{\bm{y}}-\bm{x}}{t}+\frac{\alpha}{t}(\hat{\bm{y}}-\bm{y}_{k})+\frac{\bm{x}_{k}-\hat{\bm{y}}}{t}$
$\iff\hat{\bm{y}}=\bm{y}_{k}+\frac{\bm{x}-\bm{x}_{k}}{\alpha}.$
In addition,
$\bm{x}-\hat{\bm{y}}=\frac{\bm{x}_{k}-(1-\alpha)\bm{x}}{\alpha}-\bm{y}_{k}\implies\frac{1}{2t}\left\|{\bm{x}-\hat{\bm{y}}}\right\|_{2}^{2}=\frac{1}{2t\alpha^{2}}\left\|{\bm{x}_{k}-\bm{x}+\alpha(\bm{x}-\bm{y}_{k})}\right\|_{2}^{2},$
$\bm{x}_{k}-\hat{\bm{y}}=\bm{x}_{k}-\bm{y}_{k}+\frac{\bm{x}_{k}-\bm{x}}{\alpha},$
and
[…]
$=J(\bm{y}_{k})+\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}+\frac{\bm{x}_{k}-\bm{x}}{\alpha}}\right\|_{2}^{2}+\frac{\alpha}{2t}\left\|{\frac{\bm{x}-\bm{x}_{k}}{1-c}}\right\|_{2}^{2}$
$=\langle\bm{x}_{k}-\bm{y}_{k},\bm{x}-\bm{x}_{k}\rangle/t\alpha-\frac{1}{2t\alpha^{2}}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}+\frac{\alpha}{2t\alpha^{2}}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}$
$=\langle\bm{x}_{k}-\bm{y}_{k},\bm{x}-\bm{x}_{k}\rangle/t\alpha-\frac{c}{2t\alpha^{2}}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}.$
From this, we deduce
$\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PQM}}(\bm{y})\right\}$ $=\frac{1}{2t\alpha^{2}}\left\|{\bm{x}_{k}-\bm{x}+\alpha(\bm{x}-\bm{y}_{k})}\right\|_{2}^{2}$
$\quad+\langle\bm{x}_{k}-\bm{y}_{k},\bm{x}-\bm{x}_{k}\rangle/t\alpha-\frac{c}{2t\alpha^{2}}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2},$
$=\frac{1}{2t}\left\|{\bm{x}-\bm{y}_{k}}\right\|_{2}^{2}+\frac{1}{2t\alpha^{2}}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}$
$\quad-\langle\bm{x}-\bm{x}_{k},\bm{x}-\bm{y}_{k}\rangle/t\alpha$
$\quad+\langle\bm{x}_{k}-\bm{y}_{k},\bm{x}-\bm{x}_{k}\rangle/t\alpha-\frac{c}{2t\alpha^{2}}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}$
$=\frac{1}{2t}\left\|{\bm{x}-\bm{y}_{k}}\right\|_{2}^{2}+\frac{1}{2t(\alpha}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}$
$\quad-\frac{1}{t\alpha}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2}$
$=\frac{1}{2t}\left\|{\bm{x}-\bm{y}_{k}}\right\|_{2}^{2}+\frac{1}{2t\alpha}\left\|{\bm{x}-\bm{x}_{k}}\right\|_{2}^{2},$
for some $k\in\{1,\dots,K\}$.